Help Workshop 5 Problems Student Q Search Sheet Review View * OBOU Home Insert Page Layout Formulas Data Calibri (Body) - 12 - A A
Answer #1

Working capital = Current assets - Current Liabilities

= 850,000-650,000

= 200,000

Current Ratio = Current assets/Current liabilities

= 850,000/650,000

= 1.31

Debt to assets ratio = Total Liabilities/Total assets

= 2,150,000/3,050,000

= 0.70

Debt to equity ratio = Total Liabilities /Total stockholders equity

= 2,150,000/900,000

= 2.39